### v1

# The Quantum Measurement Problem: A Review of Recent Trends

Abstract
Left on its own, a quantum state evolves deterministically under the Schrödinger Equation, forming superpositions. Upon measurement, however, a stochastic process governed by the Born rule collapses it to a single outcome. This dual evolution of quantum states—the core of the Measurement Problem—has puzzled physicists and philosophers for nearly a century. Yet, amid the cacophony of competing interpretations, the problem today is not as impenetrable as it once seemed. This paper reviews the current status of the Measurement Problem, distinguishing between what is well understood and what remains unresolved. We examine key theoretical approaches, including decoherence, many-worlds interpretation, objective collapse theories, hidden-variable theories, dualistic approaches, deterministic models, and epistemic interpretations. To make these discussions accessible to a broader audience, we also reference curated online resources that provide high-quality introductions to central concepts.
Keywords Decoherence; Measurement Problem; Quantum state collapse; Quantum mechanics interpretation
1 Introduction
Quantum mechanics, as we learn in college, poses that an isolated system evolves unitarily and deterministically following the Schrödinger Equation. When we measure the system, the Born rule determines the output probabilities. The quantum state collapses in a non-unitary, stochastic way, so the same outcome is obtained if the measurement is repeated. Despite its awkward double-evolution standard, this seemingly simple set of rules enabled comprehension of nature and technological advances to a level our species has never seen.
Nevertheless, let’s ask what “measure the system” means. We will not find an answer in the standard quantum mechanics formulation. This innocent question unravels a disturbing thread of conceptual problems, which are surprisingly difficult to address experimentally and may have dramatic philosophical consequences over our conception of the world.
The Measurement Problem has been haunting physicists and philosophers of science for almost as long as the century since quantum mechanics was proposed. We can take as an anecdotal example the World Science Festival held in the first semester of 2024, a science popularization event hosted by Brian Greene. Greene asked three quantum mechanics specialists about the Quantum Measurement Problem and got three entirely distinct answers. The entry Interpretations of Quantum Mechanics in Wikipedia contains thirteen distinct classes of theories and interpretations, and it does not even account for some recent proposals.
The cacophony of proposed solutions for the Measurement Problem is disturbing. It hurts our conception of science as delivering a consistent description of reality. Why, after a century of quantum mechanics, do we still have no consensus on its most fundamental process? What do we truly understand about the Quantum Measurement Problem?
Any newcomer trying to get an answer to these questions faces a formidable challenge: the Measurement Problem research is spread throughout uncountable isolated sub-communities. These groups don’t dialogue with each other, making it exceptionally difficult to build an overview of what is going on. The large number of single-author papers doesn’t help either. These papers raise hypotheses and propose theories and tests. However, most don’t resonate within a mature community and only add to the cacophony.
As newcomers in the field ourselves—we just published our first contribution recently [1]—we felt firsthand all these issues. We tackled this challenge with extensive reading, discussion, and consult with specialists on specific topics. We explored all corners of the field rather than focusing on a single subfield. This review arose as a natural consequence. Our goal is to deliver a broad overview of the several competing research lines. We don’t aim at a comprehensive survey (which should not be possible within reasonable limits of a paper) but to illuminate the main questions being currently discussed, allowing curious non-specialists to start digging the literature by themselves. In doing so, we hope to show that the Measurement Problem is not as bleak as it seems, although there is still much to uncover.
Given the Measurement Problem’s central role, it’s unsurprising that several other reviews have been recently published to explore its unresolved aspects from different perspectives. Hance and Hossenfelder [2] critically assess whether the problem is fundamental or an artifact of our theoretical framework. In a synthetic review, they identify five requirements that any satisfactory solution to the problem must satisfy and analyze several proposed solutions. Müller [3] followed a different path. Instead of focusing on the proposed solutions, he approaches the measurement problem from a formal logic perspective. He decomposes it into six independent problems, which should be investigated separately. Bassi, Dorato, and Ulbricht [4] provide an in-depth analysis of objective collapse models, discussing them from theoretical, experimental, and philosophical perspectives. It perfectly updates Bassi and co-workers’ excellent (and extensive) review published in 2013 [5]. Meanwhile, Freire Jr. [6] delivers a historical account, linking the foundations of quantum mechanics to current scientific and technological advances.
Each of these works is worth the attention of readers interested in the status of foundational discussions of quantum mechanics. Nevertheless, none deliver a broad coverage of the Measurement Problem; that’s precisely what we missed when we started our journey. Addressing this gap, our review organizes the various proposed solutions into a structured overview, offering a broad reference point for both specialists and newcomers.
Our work also brings a distinct perspective by examining the Measurement Problem through the chemistry lens (since two authors, RSM and MB, are theoretical chemists). Historically, chemists have treated quantum mechanics pragmatically, using it as a tool rather than engaging with its foundational questions. However, recent advances in quantum materials [7], quantum dots for imaging and drug delivery [8], ultrafast spectroscopy [9], and quantum computing [10, 11] have brought the Quantum Measurement Problem into sharper focus at the chemistry labs. These fields increasingly rely on quantum coherence, entanglement, and decoherence—concepts central to quantum measurement theory [12]. As chemists push the boundaries of quantum state control, they may need to reconsider long-standing assumptions about wave function collapse, observer effects, and the role of decoherence in chemical processes.
Despite being unconventional in scientific papers, we extensively referred to non-peer-reviewed web resources, including books, essays, and videos. These resources offer plain language introduction and analysis without intimidating technicalities. They naturally don’t replace technical literature but should not be neglected as relevant entry points to the subject.
2 Quantum State Description and Evolution
Quantum mechanics encodes all information about a physical system in the quantum state , a complex-valued vector in the Hilbert Space containing contributions from all potential measurement outcomes.
Consider, for example, the ammonia molecule NH3 (Figure 1). Its more stable geometry resembles a pyramid, with the three hydrogen atoms forming a triangular basis and the nitrogen atom lying above or below it. Thus, ignoring molecular vibrations and rotations, the nuclear quantum state of ammonia has two components, corresponding to pyramids up and down [13, Chs 8 & 9]. Let’s name them and states, respectively, which means that we will treat NH3 as a quantum bit (or a qubit).
In this example, the general quantum state is a bi-dimensional vector (see Figure 2) written as
where and are complex-valued coefficients (or amplitudes) with .
Eq.(1) trivially generalizes to any number of dimensions as . For a continuous basis, it becomes , where the amplitudes are called wave functions.
is usually a dynamic quantity evolving according to the Schrödinger Equation
where is the Hamiltonian operator. Alternatively, the dynamics of can be described in terms of the time evolution of the amplitudes, as indicated in Figure 2(c).
When we measure the system, we always observe a single outcome, an eigenvalue of the Hermitian operator representing the observable. The probability of each outcome is the square of the modulus of the corresponding state amplitude at the measurement time . That means,
When we determine the outcome value, we say the quantum state collapsed.
Alternatively, we can describe the collapse in terms of projector operators. In our example, the operators are
Suppose that the molecule is initially in state given by Eq. (2) at the measurement time . If we project on state , we have the trivial result
Thus, the probability of getting is
and the state after measurement is
Naturally, similar equations hold for measuring . Eqs.(6) and (7) are entirely equivalent to Eq.(3) (apart from a global phase) but only refer to the projection operator, not the amplitudes. The advantage of this formalism, called projection-valued measure, is that it describes the collapse that can be generalized to ensembles of states [14].
All this discussion about collapse has assumed that it is instantaneous and occurs at tm. Standard quantum theory gives no clue on whether this is precise or not. Recent resonant inelastic X-ray scattering (RIXS) experiments have suggested that state collapse occurs gradually over a finite time [15], but this interpretation remains ambiguous, as the observed dynamics could also be explained by progressive decoherence rather than a fundamental modification of quantum measurement. We will return to when and how long collapse takes place in the discussion of Objective Collapse Theories in Section 5.3. The implications of an instantaneous collapse are explored in Section 5.7, in the context of quantum field measurements.
3 The Problems with Quantum Measurements
To summarize the description of the quantum state evolution discussed in the previous section, standard quantum mechanics prescribes that a quantum state follows two types of time evolution: Left on itself, it evolves with the Schrödinger Equation, which is unitary and deterministic [16, Ch 3][17, Ch 1]. During a measurement, however, the quantum state evolves with the Born rule, which is non-unitary and stochastic.
There are different ways to scrutinize this description [3, 18, 19]. Here, we follow Schlosshauer [20, Ch 2], who decomposed it into three problems: the problem of the preferred basis, the problem of the nonobservability of interference, and the problem of outcomes. Each of these challenges arises naturally from the standard formulation of quantum mechanics and highlights a gap in our understanding of what happens when a quantum system is measured. Below, we define these problems in sequence.
The problem of the preferred basis: Suppose we have an apparatus to detect whether the NH3 molecule is up or down. This apparatus is also supposed to be a quantum system, which is initially in the state “ready.” If it measures the molecule up, its pointer moves to 1. If it detects down, it moves to 0:
If the molecule is initially in the superposition
the interaction with the apparatus will lead to the entangled state
Note that we assumed the interaction does not alter the system’s state. This one-to-one correlation between the system and the measurement apparatus establishes the von Neumann measurement scheme [20], which now only depends on quantum state collapse to determine the outcome.
We are free to change the basis representation in quantum mechanics. The states of the molecule can also be described, for example, on the following basis (Figure 3):
which has a one-to-one association with the apparatus basis:
In this new basis, the premeasured system-apparatus state is
Thus, it would appear that once the apparatus measured the molecule in the basis, the measurement in the (or any other arbitrary basis) is automatically determined. This contradicts the fundamental principle of quantum mechanics, as the observables related to and are non-commutative (note that the previous relations are equivalent to the Pauli matrices and ) and cannot be simultaneously well-defined. Moreover, this description is also unsettling, as the apparatus set to measure is physically distinct from a setup to measure .
We know that in practice, the apparatus designed to measure will produce outputs on this basis. Therefore, we must explain what eliminates basis ambiguity in quantum mechanics formalism, selecting a specific basis. This is the problem of the preferred basis, first spelled out by Zurek [21].
The problem of the nonobservability of interferences: Quantum mechanics predicts that systems can exist in coherent superpositions of states, and interference effects should be observable, such as interference patterns in double-slit experiments. However, in everyday macroscopic systems, we do not see such interference. Instead, objects appear to have well-defined classical properties. Why do we not see superpositions everywhere?
The ammonia molecule from our example may, in principle, exist as a superposition of its ’up’ and ’down’ states. Still, measurements always find it in one or the other, never in a visibly interfering combination. We may feel tempted to attribute it to tiny de Broglie wavelengths of large systems. However, this limitation can be circumvented, for instance, by using effects such as Talbot diffraction [22]. Under carefully controlled experimental conditions, interference patterns have been reported for molecules as large as 2000 atoms [23]. However, it requires isolation from environmental disturbances never observed in ordinary mesoscopic systems. This raises the fundamental question of how the environment suppresses interferences.
The problem of outcomes: Even after addressing the previous two issues—identifying a preferred basis and suppressing interference—there remains a fundamental question: why do measurements yield definite outcomes rather than superpositions? The final state of the system-apparatus interaction is still a quantum superposition of different possible results, yet we never observe such superposition.
Furthermore, even after a basis is selected and interference is suppressed, quantum mechanics does not explain why a specific outcome occurs in each measurement. The Born rule successfully predicts the statistical distribution of outcomes over many repeated experiments, but it does not explain why a particular result is obtained in a single instance.
4 Decoherence
The problem of the nonobservability of interferences is addressed by decoherence. Decoherence disperses the quantum information about a system into correlations with the surrounding environment, suppressing the observable effects of superposition [20, 24]. To have a feeling of how decoherence acts, suppose that our and system can interact with an environment with only three states, , , and . We initially prepare the system in a superposition of and and the environment in , such that the initial state of the system plus environment is
Because of the system-environment interaction, the quantum state evolves into an entanglement:
We can monitor the system’s evolution alone, averaging it over the environment states. This is done by computing the reduced density matrix for the system. This average is obtained by taking the trace over the environment’s states of the full-density operator:
The diagonal terms and are the and populations. The off-diagonal terms are the coherences and are responsible for quantum interferences.
In general, the environmental states are not orthogonal, and the coherences are non-null. However, as time passes, the environment states tend to become orthogonal because their dynamics are different when the environment interacts with the system’s and geometries. Thus, the off-diagonal terms tend to zero
because dynamically becomes null. This process is called decoherence.
In formal terms, decoherence arises from the time propagation of the reduced density matrix, which evolves with a master equation of the type [20, Ch 4].
where is the system’s Hamiltonian, including the environment’s perturbation, and is the term responsible for decoherence to the environment.
Decoherence tends to be extremely fast when considering the environment as a multidimensional system. For example, we can employ the Joos-Zeh model [25] to estimate the decoherence time of the and superposition in ammonia gas. According to this model, the coherence in Eq.(17) decays exponentially
with a decoherence time given as
In this equation, is the number density of the gas at temperature . is the molecular mass. is the molecule’s size (the molecule is assumed to be a dielectric sphere), and is the displacement between the and geometries. Plugging these quantities in Eq.(17), we find that NH3 coherence disappears almost instantaneously within seconds under room temperature and pressure. Although the superposition persists in the combined system and environment, its effects become unobservable due to dispersal in the gas.
Although decoherence typically occurs in femtosecond or even sub-femtosecond timescales, specific molecular systems exhibit surprisingly long-lived coherence effects. Kaufman et al. [26], for instance, demonstrated that electronic coherences could survive for hundreds of femtoseconds in molecules with parallel potential energy surfaces due to reduced dephasing between branched wavepackets (see also 4). Schultz et al. [12] provide a comprehensive review of cases where vibronic coupling and structured environments lead to long coherence lifetimes, including examples from light-harvesting complexes in photosynthesis.
Babcock et al. [27] reported superradiant effects in tryptophan networks as large as microtubules, which require long-lived coherent effects. Their result is astonishing, given that fluorescence requires nanoseconds, and it isn’t easy to rationalize how coherence could survive for such a long time. For comparison, Tegmark estimated the decoherence time in microtubules to be shorter than 10-13 s [28]. Babcock’s result has also attracted media attention (see here and here) because it may support Hammeroff-Penrose’s hypothesis on the quantum origin of consciousness (see Section 5.6.2).
These findings suggest that molecular architecture can sometimes counteract decoherence, challenging the assumption that coherence always vanishes rapidly. Theory, especially atomistic simulations [29], has still to catch up in this field.
In addition to suppressing quantum interferences, decoherence also addresses the problem of the preferred basis. It leads to the selection of a stable outcome basis through a process known as einselection, which drives the quantum state to the basis least entangled with the environment [21]. This preferred basis depends on the Hamiltonian [30]. Assuming that the position operator commutes with the system-environment interaction Hamiltonian (which is usually the case), decoherence tends to select the position eigenstates as the preferred basis as long as the interaction Hamiltonian is more important for the dynamics than the system’s Hamiltonian. If both terms compete, the preferred basis is close to coherent states, which minimizes the position and momentum uncertainty. If the system’s Hamiltonian dominates the dynamics, energy eigenstates are preferred as the basis.
For a molecule in a gas, Paz and Zurek analysis [30] implies that the outcome basis the environment selects for measuring nuclei is approximately position states. On the other hand, energy states emerge as the preferred outcome basis for measuring electrons. Thus, when we initially chose the basis for describing the quantum state of the ammonia molecule, we did it because it matched our classical intuition of nuclei corresponding to particles with defined positions. After discussing decoherence, we realize that we may have developed such a classic intuition because the environment selects the basis and suppresses superposition effects, creating the world of particles with well-defined positions.
Nevertheless, einselection may not be the last word in the preferred basis problem. Adil and -coauthors have recently shown that the preferred basis is not solely dictated by the system-environment interaction but also by the natural factorization of the global Hilbert space [31]. Different factorizations can lead to multiple coexisting sets of pointer states, meaning classical behavior emerges in different ways depending on the structure of the Hamiltonian. These co-existing classical realms challenge the idea that einselection always yields a single classical outcome. They may even reignite the preferred basis problem, but now at the level of Hamiltonian factorization.
Decoherence is part of the standard quantum mechanics toolkit. It is an experimentally well-documented effect crucial for quantum mechanical applications, such as quantum information [14]. Regardless of one’s preferred interpretation, the role of decoherence remains essential and cannot be ignored.
This dominant role of decoherence in the current understanding of quantum mechanical processes doesn’t imply it is a consensus. Landsmann, for instance, characterizes decoherence as “an unmitigated disaster” [32, Ch 11]. His core criticism is that decoherence rigorously only occurs in an environment with infinite degrees of freedom after infinite interaction time with the system. Kastner reinforces this criticism by noticing that decoherence claims of describing irreversible processes suffer from similar conceptual problems already present in statistical mechanics when explaining the emergence of irreversibility from reversible processes [33].
5 Facing the Problem of Outcomes
In the previous section, we have seen that decoherence clarifies a significant portion of the Measurement Problem. However, the problem of outcomes is still open: After the basis is chosen and quantum superpositions are suppressed, the system still remains on a mixture of possible outcomes. Decoherence does not tell how and why only one of these outcomes is measured. At this point, standard quantum mechanics interpretation claims the occurrence of an additional dynamic step, the collapse:
We can illustrate this point with an example of a photoexcited molecule undergoing internal conversion Figure 4. Light initially promotes the molecule to an excited electronic state. The nuclear wavepacket evolves on the potential energy surface of the excited state until it reaches a degeneracy region with the ground state potential energy surface. There, quantum coherence builds between the two states, creating a state superposition. The dephasing between the wavepacket components in each state counteracts, causing decoherence. All events up to this point, including decoherence, can be simulated by propagating the Schrödinger equation for the molecule [34]. If the molecule is measured, it will either be in the ground or excited state. However, our simulation can only tell the probabilities of these outputs.
Standard quantum mechanics assumes a wave function collapse occurred. Nevertheless, the existence of collapse is far from consensual. It is regarded as a superfluous hypothesis [35] in several quantum mechanical interpretations.
In the following subsections, we will quickly dive into a few classes of potential solutions that represent the current debate on the Measurement Problem, with little attention to philosophical and historical aspects. More comprehensive approaches can be found in Refs.[6], [32, Introduction & Ch 11], [36, 37].
Before that, we should clarify the concepts we are dealing with. In literature, it is common to use decoherence and collapse concepts interchangeably. We should avoid it, as these phenomena may, in principle, be completely independent and occur in different time scales.
As stated in Eq. (21), decoherence acts on the off-diagonal terms of the system’s reduced density matrix. Collapse, however, zeroes the diagonal terms except for one element. We can be a bit more flexible with this definition. In principle, a quantum theory including collapse could go as far as to provide a diagonal density matrix. This means we would be left in a situation of statistical uncertainty: the collapse occurred, but the theory only informed us of the probability of each outcome. We could call such a case an epistemic collapse. If, however, the theory zeroes the diagonal (except one) elements, then we would have an ontic collapse.
5.1 Many-Worlds Interpretation
The lack of information on the outcomes is not a problem for some interpretations. For instance, the Many-Worlds Interpretation (or Relative State Interpretation or, still, Everettian Quantum Mechanics) proposes that the remaining outcomes after decoherence correspond to independent universe branches [38]. Observers existing in each branch perceive different outcomes because their quantum state, entangled with the molecular quantum state, also branches. In one branch, the combined molecule-observer quantum state may be and in another, . There are no branches with superpositions of both states because of decoherence toward the environment surrounding the system and the observer suppressing them [39]. Thus, the Many-Worlds Interpretation posits the existence of a single quantum universe composed of non-communicating classical branches.
For the Many-Worlds Interpretation, the quantum state is a complete description of reality (no hidden variables are required), exclusively evolving according to the Schrödinger Equation (collapse is only an illusion caused by the universe branching) [40, 41]. Thus, the Many-Worlds Interpretation is consistent with standard quantum mechanics and requires no theory modification. However, it remains an interpretation rather than a theory until it gains experimental testability. A significant challenge is the technical infeasibility of maintaining and measuring a quantum system in a coherent superposition of orthogonal states, especially for large systems. This difficulty explains why superpositions of macroscopic states, such as and , cannot be observed directly due to rapid decoherence. Susskind and Aaronson demonstrated this limitation using methods from quantum computational complexity, naming it the quantum-hard problem [42, 43].
A puzzling feature of the Many-Worlds interpretation is understanding how the deterministic branching relates to the wave function amplitudes and the Born rule. Carroll and Sebens [44] propose that this relation is established after decoherence occurs but before the measurement is registered. In this period, if the observer wants to estimate which branch they are, their rational credence directly leads to the Born rule. Thus, the probabilistic character of quantum mechanics would be a subjective aspect arising from the observer’s self-locating uncertainty. Carroll and Sebens’ derivation of the Born rule is only strictly valid if the ratios of branch amplitudes are the square roots of rational numbers. However, they argue that the large number of branches in practical situations always allows this condition to be approximately satisfied. This approximation also allows for experimental tests of their hypothesis by searching for small deviations in the Born rule. Kent [45] forcefully criticized Carroll and Sebens’ argument regarding their conceptual foundation and mathematical rigor. His main criticism is that an underlying ontology supporting Carroll-Sebens’ model would require a unique and objective branching scheme.
Like Carroll and Sebens, Short [46] derives the Born rule in Many-Worlds Interpretation based on estimating the probability of picking a random branch among many. However, instead of relying on self-locating uncertainty, he does it by establishing three axioms that the probability must satisfy: 1) the probability depends only on the present state (regardless of how it was generated); 2) only branches with non-zero amplitudes count as components of the many-world state; 3) probability cannot flow between uncoupled worlds. Short then shows that the Born rule uniquely satisfies these axioms.
The Many-Worlds Interpretation may seem exotic to non-specialists due to its exuberant ontology. Despite this perception, it is a serious contender among the proposed solutions to the Measurement Problem. It is a popular interpretation among physicists, with prominent supporters like Schlosshauer [20], Carroll [40, 41, 44], Zurek [39], Zeh [35], and Deutch promoting different variants. However, it also faces outstanding critics, like Penrose [47] and Albert. For Landsmann, the Many-Worlds Interpretation would be “acceptable only if truly everything else has failed” [32, Ch 11].
5.2 Epistemic Interpretations
Epistemic interpretations claim that a quantum state is no more than a mathematical tool to predict outcomes. This is the stance of interpretations such as the Relational Interpretation [48, 49, 50], Quantum Information-Theoretic approaches [51], Quantum Bayesianism [52], and Indivisible Stochastic Process Theory [53], which are discussed next.
5.2.1 Relational Interpretation
The Relational Interpretation of quantum mechanics (or relational quantum mechanics), proposed by Rovelli, focuses on the interaction between systems, emphasizing that properties are defined only in relation to such interactions [50]. Thus, the molecular geometry may be well-defined for a nearby molecule in the environment acting as an observer while remaining in superposition for another molecule that is too distant to interact significantly.
Rovelli draws attention to the fact that this state dependence on a specific observer is not new in physics. It is analogous, for instance, to a velocity dependence on the particular reference frame where it’s measured. The novelty in quantum mechanics is that the non-commutative nature of quantum variables implies that not all properties can be sharply defined simultaneously [50]. This limits the completeness of information derivable from a single interaction.
In the Relational Interpretation, the wave function is not ontologically real (see Section 6 for discussion). Instead, it serves as a tool for predicting the probabilities of interaction outcomes. Conceptual issues with measurement arise only when undue ontological weight is assigned to it. By treating the wave function purely as a predictive instrument, measurement can be reframed as an ordinary interaction, resolving these difficulties.
The Relational Interpretation is characterized by a sparse ontology [50], meaning that properties are not defined between interactions. Nevertheless, it remains a realist interpretation because it posits a physical world of interacting systems.
Giacomini, Castro-Ruiz, and Brukner have recently extended relational ideas beyond Rovelli’s Relational Quantum Mechanics by incorporating Quantum Reference Frames, QRF [54]. In their approach, a QRF is simply a quantum system elevated to the role of a reference frame so that its degrees of freedom define the origin relative to which one describes other systems. By extending canonical transformations to the quantum domain, Giacomini et al. show how usual frame changes (translations, for example) remain valid even if the reference frame itself is in a superposition or entangled state. This leads to the relativity of notions like superposition and entanglement. If a given particle is in a spatial superposition from the lab’s viewpoint, then from the particle’s viewpoint, the lab may itself appear in a corresponding superposition.
This perspective may offer new insights into the Measurement problem, particularly in Wigner’s friend scenarios, where the possibility of conflicting accounts among observers becomes natural if each observer constitutes its own QFR [54, 55]. In Section 5.3.2, we discuss how Giacomini and Brukner use QFRs to challenge Pennrose’s conjecture on gravity-induced collapse [56].
5.2.2 Information-Theoretic and QBism Views
Another epistemic perspective comes from Quantum Information Theory, which treats quantum mechanics as an axiomatic framework akin to thermodynamics. Brukner and Zeilinger [51] argue that quantum mechanics should be understood as a theory of information rather than a theory of reality, where the fundamental postulates describe how information is processed rather than how physical systems evolve.
Recent experimental results reinforce this perspective. Spegel-Lexne and co-authors experimentally demonstrated the equivalence of entropic uncertainty and wave-particle duality [57], reinforcing that quantum measurement can be understood through informational constraints rather than wave function collapse.
The epistemic perspective in Quantum Information Theory aligns with Quantum Bayesianism (QBism), which takes a radically subjective approach: the quantum state does not represent an objective reality but rather an observer’s personal beliefs about measurement outcomes [52]. When a measurement is made—such as determining if the molecule is in the state—the observer updates the quantum state to reflect the new information [52, 58].
However, Information-Theoretic interpretations, particularly QBism, face several challenges. The Pusey-Barret-Rudolph (PBR) theorem [59] states that epistemic interpretations of the quantum states cannot entirely reproduce the results of quantum mechanics. It has also been claimed that epistemic interpretations (at least in the QBism form) contradict Ozawa’s Intersubjectivity theorem, which poses that two observers jointly measuring the same system must obtain the same outcome [60, 61]. However, no-go theorems can always be criticized for their underlying hypotheses [62] (We come back to this point in Section 5.4.).
5.2.3 Indivisible Stochastic Process Theory
Barandes has recently proposed the Indivisible Stochastic Processes theory, where classical configurations of particles, fields, or qubits evolve via non-Markovian dynamics [53]. In this type of dynamics, which recovers all results from conventional quantum mechanics, transition probabilities cannot be estimated at intermediate times. In Barandes’ approach, wave functions are considered predictive tools only, and the Hilbert space of quantum mechanics is demoted to a convenient instrumentalist framework without an ontological basis.
Barandes draws a historical analogy, comparing the transition from instrumentalist frameworks in quantum mechanics to the Copernican revolution, where the instrumental Ptolemaic system of epicycles was replaced by a more fundamental description based on forces. Similarly, he advocates for a clear physical picture to supersede the abstract formalism of Hilbert space mechanics. Like the Relational Interpretation (Section 5.2.1), the Indivisible Stochastic Process formulation features a sparse ontology, emphasizing that properties are not well-defined between interactions but are instead contextualized through the system’s stochastic dynamics. By embedding quantum systems within generalized stochastic systems, this approach aligns with a realist philosophy, reinterpreting quantum systems as governed by dynamic interactions without requiring measurement to be treated as a distinct or special process.
5.3 Objective Collapse Theories
5.3.1 Agnostic Objective Collapse
Diametraly opposed to Many-Worlds and Epistemic Interpretations, Objective Collapse Theories consider that collapse is a physical process, and that standard quantum mechanics is still incomplete [4, 5]. Thus, these theories attempt to change the Schrödinger Equation to account for the collapse. This subsection deals with agnostic objective collapse, that is, theories that don’t attribute a cause to the collapse. In the following subsection, we will look at gravity-induced collapse.
Most objective collapse theories predict that a system in quantum superposition randomly localizes in one of the outcomes given enough time [5, 63, 64, 65]. The quantum state is driven by a Stochastic Schrödinger Equation, such as the Gisin Equation [66, 67]
In addition to the usual Hamiltonian term, the right-hand side contains two additional terms that tend to localize the state in one of the eigenstates of the operators , as illustrated in Figure 5 for . The collapse is stochastic and controlled by a Wiener process (with and ), which recovers the Born rule. A positive real constant tells how wide the noise is and how fast the collapse is. The equation is nonlinear due to the terms. However, the evolution of the ensemble density is linear thanks to the quadratic last term [67]. This linearity is required to avoid superluminal information exchange.
Different objective collapse theories have been proposed, corresponding to specific choices of operators , including the Hamiltonian operator [68], position operators (GRW after the initials of the authors [69]) density number operators (CSL for Continuous Spontaneous Localization [64]), and mass density operators [70] (which is discussed in Section 5.3.2). These specific models are thoroughly reviewed in Ref. [4].
5.3.2 Gravity-Induced Objective Collapse
The highest stakes among the objective collapse theories are those proposing that gravity is the cause of the collapse. Penrose [71], for instance, claims that distinct mass distributions in superposition ( for up or for down in ammonia molecule) would cause slightly different spacetime distortions. The instability associated with these superimposed distortions leads to the wave function collapse within a time
where is the Newtonian gravitational self-energy
and is the gravitational constant. This collapse time can also be derived by computing the difference between Newtonian free fall accelerations in the space surrounding each geometry [72]. A direct application of Eq. (23) predicts it would take billions of years to collapse the wave function of an isolated molecule, but only seconds to observe collapse in a kg body (Figure 6) [1].
The same collapse time arises from the Quantum Mechanics with Universal Density Localization (QMUDL), the model Diósi proposed for the stochastic state evolution [70]. In QMUDL, the Schrödinger Equation is modified to
where the scalar Wiener process is defined by and . This equation is a particular case of Eq.(22) where the operators are taken proportional to the mass density operator . The random noise is proportional to the ratio between the gravitational constant and the reduced Planck constant . is an arbitrary dimensionless constant. Despite the distinct and independent derivations, the collapse time (23) is generally called the Diósi-Penrose Model.
The Diósi-Penrose model has attracted criticism on multiple fronts. Gao [73] raised several issues on the Penrose conjecture. His main point is that Penrose’s argument by analogy with the conventional quantum mechanical time-uncertainty principle is not rigorous enough. Moreover, he is also right to point out that although we commonly refer to Diósi-Penrose model, Diósi and Penrose theories are fundamentally distinct, only sharing, maybe by simple coincidence, the same collapse time. Giacomini and Bruckner [56] also challenged Penrose’s conjecture that quantum superposition conflicts with Einstein’s equivalence principle. Using the Quantum Reference Frames (QRF) formalism (see also 5.2.1), they showed that it’s possible to generalize the principle (or, more precisely, always to make the metric field locally Minkowskian) even in the presence of quantum superpositions of massive bodies.
Additionally, we performed atomistic simulations to show that the Diósi-Penrose collapse time reaches a minimum value independent of superposition displacement and, in some instances, can be surprisingly long [1]. Figurato et al. [74] has also independently concluded that the model could have too long collapse times for macroscopic bodies. Beyond these theoretical concerns, experimental tests have also challenged the model. Donadi et al. [75] and Arnquist et al. [76] conducted an underground experiment to test gravity-related collapse mechanisms. Their results placed stringent constraints on the model’s parameters.
Although all these points must be considered when assessing the validity of the Diósi-Penrose model, none of them directly falsify the model. And, as Thorne said, ”Penrose has a way of always being right, even when he does things in a strange way” [77].
Oppenheim has recently proposed another hypothesis that attributes wave function collapse to the influence of gravity. Indeed, his Postquantum Theory of Classical Gravity does not focus on the measurement problem and has a more ambitious goal of reconciling gravity and quantum mechanics [78]. Oppenheim’s central hypothesis is that gravity is a classical field. To ensure classical gravity can interact coherently with quantum matter fields, Oppenheim developed a Classical Quantum (CQ) Dynamics framework, which circumvents the usual obstacles encountered by semiclassical models, such as violation of the superposition principle due to nonlinearities in the Schrödinger equation [78, 79, 80].
The CQ dynamics of the hybrid state (where are the classical phase space coordinates) is expressed in terms of the general master equation [78]
which can describe back reactions of quantum matter fields in classical (including gravitational) fields. This equation has completely positive dynamics (hence probabilities remain positive), preserves the trace (probabilities are conserved), and is linear in the density operator (superluminal signaling is avoided). In addition to the commutator term responsible for the unitary evolution and the Lindbladian term [81] (proportional to ) causing decoherence, the integral term promotes jumps in the quantum and classical subsystems. The last term preserves the norm of the quantum state.
Like in modeling quantum open systems [82], Eq.(5.3.2) is not directly solved for but rather through an ensemble of stochastic trajectories (unraveling procedure) [83, 84]. In each trajectory, the quantum degrees of freedom undergo stochastic projection onto the eigenstates of the corresponding Lindblad operators, collapsing to a definite outcome in a manner that statistically recovers the Born rule. These quantum jumps induce corresponding phase space jumps in the classical degrees of freedom, representing a backreaction of the quantum system on the classical one. In addition to collapse, this equation also describes decoherence, leading to the suppression of quantum coherence over time. Figure 7 illustrates the quantum and classical stochastic jumps in a system composed of a qubit and a classical particle discussed in Ref.[83]
Even though stochastic trajectories are introduced as a means to propagate the density evolution (5.3.2), Oppenheim and coauthors assign them ontological significance [83]. A stochastic trajectory represents the system collapsing to a specific state, just like it is also done in the other objective collapse theories that we discussed above. Oppenheim’s theory implies that collapse does not require measurement, as it occurs naturally due to interactions with a classical field [83].
Gravity-induced wave function collapse may have profound implications for physics. If any of these theories are confirmed, it would not only close the Measurement Problem but also imply that gravity plays a unique role in nature, different from the other fundamental interactions. Moreover, it would also impact future efforts to reconcile general gravity and quantum mechanics, moving toward what Penrose calls gravitization of quantum mechanics [85].
5.4 Hidden-Variables Theories
5.4.1 Physical Restrictions on Hidden Variables
Hidden variable theories (also known as ontic theories) explore the hypothesis that the quantum state may not be a complete description of the system and that a set of additional (hidden) variables may contain further information (accessible or not) [86]. The hidden variables discussion is where the Measurement Problem spills the most on other fundamental questions of quantum mechanics, like nonlocality in entangled systems, the completeness of quantum mechanics, and the existence of properties independently of measurements.
The hidden variable theories are usually framed within the limits of two no-go theorems and a couple of physically motivated restrictions. The theorems are the Bell and the Kochen-Specker ones. The Bell’s theorem states that no local hidden-variable theory can be entirely compatible with quantum mechanics [87, Ch 2] [88]. Local in this context means that “the result of a measurement on one system be unaffected by operations on a distant system with which it has interacted in the past” [87, Ch 2].
In turn, the Kocher-Specker theorem [88, 89, 90] states that any hidden variable theory consistent with quantum mechanics must renounce either all observables having definite values at all times or these values being independent of how they are measured. The theorem is strictly valid for Hilbert spaces with dimension superior to two.
Finally, the physical restrictions are the non-signaling principle (imposing that nonlocality cannot be harvested to exchange superluminal information) and the statistical independence (supposing that initial statistical setups can be completely uncorrelated; see also Section 5.5).
The situation seems simple from there. For instance, the validity of Bell’s Theorem–which experiments have vastly supported [91]–seems to imply that physical theories must either be non-local or not have properties defined at all times. However, as mentioned before, no-go theorems are as strong as their underlying hypotheses, often deeply hidden in the deduction. The restrictions imposed by these theorems and physical principles have been constantly challenged [92, 93] (except for the non-signaling principle, which is likely the only consensual rule in the game). Relaxing statistical independence (sometimes loosely called the free-will principle) in Bell’s theorem, for example, may allow for building consistent local and realist theories. We will return to this point in Section 5.5.
5.4.2 Bohmian Mechanics
Most of the practical development in hidden-variable theories happens in the frame of Bohmian Mechanics, which poses the existence of corpuscules moving in a non-Newtonian way and guided by a wave function [94]. The corpuscule’s velocities are determined by a velocity field (the de Broglie-Bohm Equation) dependent on the wave function, which, in turn, follows the Schrödinger evolution. This deterministic time evolution in Bohmian Mechanics automatically solves the problem of outcomes. Although Bohmian Mechanics’ experimental predictions are equivalent to quantum mechanics [95], it is not discarded that they could be experimentally tested [96].
Poirier has proposed an alternative theory prescribing non-Newtonian trajectories but entirely discarding wave functions [97, 98]. In this case, the quantum system is described by an ensemble of interacting deterministic real-valued trajectories with well-defined positions and momenta. This type of theory inscribes itself into the tradition of hydrodynamic interpretation of quantum mechanics, inaugurated by Madelung right after Schrödinger’s original work [99][100, p. 222]. In Madelung’s interpretation, quantum mechanics is modeled as a quantum fluid [101], which can be understood as many interacting classical worlds [102].
This picture of many interacting classical worlds has been the basis of molecular dynamics simulations, where the nuclear wave function time evolution is commonly approximated by interacting and non-interacting trajectories, depending on the approach [103, 104]. Tully, who proposed surface hopping, the most popular method for nonadiabatic dynamics simulations, employed the Madelung Hydrodynamic Interpretation to justify using mixed quantum-classical approximations [105]. These molecular dynamics simulations, however, adopt the trajectory approach pragmatically without any foundational claim.
5.5 Deterministic Models
Bell’s theorem shows no local hidden variable theories can reproduce all quantum mechanical predictions [87]. Thus, the experimental violation of Bell’s inequalities seemed to be the end of any local deterministic model. However, Bell’s theorem explicitly requires the validity of statistical independence assumption, which states that the experimental settings are not influenced by any hidden factors determining the system’s evolution. This section surveys two local deterministic models—Superdeterminism and Retrocausality—built on the challenge of the statistical independence assumption.
Superdeterminism posits that hidden correlations between measurement settings and system properties predetermine quantum outcomes [106]. For instance, a recent proposal by Donadi and Hossenfelder [107] introduced a local, deterministic model for wave function collapse that explicitly violates statistical independence while reproducing standard quantum predictions. Their model suggests that all quantum states, except measurement eigenstates, are unstable under hidden-variable perturbations and deterministically collapse before reaching the detector. Unlike objective collapse models (see Section 5.3), the collapse dynamics in the Donadi-Hossenfelder model are entirely deterministic.
Recent Cosmic Bell Experiments have provided stringent tests of the statistical independence assumption in Bell’s theorem. These experiments avoid the possibility that hidden variables could influence detector settings by randomly determining measurement choices using distant astrophysical sources. By doing so, they attempt to close the freedom-of-choice loophole, which is crucial for evaluating the plausibility of superdeterministic explanations. In a landmark experiment, Rauch et al. [91] used high-redshift quasars—whose light originated billions of years ago—to set measurement parameters in a Bell test. Their results remained consistent with standard quantum mechanics, implying that superdeterministic correlations, if they exist, should date back to the early universe. Similarly, Handsteiner et al. [108] employed light from distant galaxies to eliminate potential local influences on detector choices. While these results do not strictly rule out superdeterminism, to remain viable, it must account for cosmological-scale correlations, raising deep challenges for its plausibility.
Retrocausal Interpretations also challenge the statistical independence assumption, but, even bolder than Superdeterminism, it assumes influence from future events. Such interpretations propose that the Measurement Problem arises from assuming a one-way flow of time. Models such as the two-state vector formalism [109] describe quantum systems by an initial state evolving forward in time and a final state evolving backward from the measurement outcome. The interaction of these forward- and backward-evolving states determines the probabilities of different results. The Pechukas’ forces [110], which are still the main justification used today for momentum rescaling along nonadiabatic coupling vectors in molecular dynamics [111], were derived from this same type of forward and backward propagation to determine quantum transitions in atomic collisions.
Retrocausality is still an active research field. Adlam [112] proposed that accepting nonlocality and imposing relativistic constraints (such as the nonexistence of a preferred reference frame) leads to retrocausality. Sutherland [113] derives the Born rule within a retrocausal model. This is a significant result because it changes the status of the rule from a postulate to a natural consequence of incorporating time-symmetric boundary conditions in quantum mechanics. A notable recent development is Ridley and Adlam’s Fixed-Point Formulation [114], which presents an atemporal, all-at-once framework for retrocausality.
Both Superdeterminism and Retrocausality remain controversial and experimentally unverified. Nevertheless, despite their speculative nature, these models provide a striking contrast to interpretations that rely on wave function collapse, branching universes, or fundamental randomness [115], keeping the possibility of a fully local and deterministic resolution to the Measurement Problem open.
5.6 Dualist Collapse Hypotheses
5.6.1 Classical-Apparatus Inducing Collapse
The core of the Copenhagen Interpretation of quantum mechanics, developed by Bohr and Heisenberg, is the assumption that any experiment in physics must be described in terms of classical physical concepts [32, Introduction]. When such an experiment is in the realm of quantum mechanics, the tension between the system’s quantum description and the apparatus’s classical description is the crucial factor causing statistical uncertainty and ultimately selecting a single output [116]. There is no claim that the classical apparatus is not composed of quantum particles. It is just that the description of the experiment is restricted to the use of classical terminology.
This quantum-classical duality, formulated as the Heisenberg Cut arbitrarily separating classical from quantum systems, still plays a role in a pragmatic definition of measurement [117]. It is also at the basis of the asymptotic emergence hypothesis, which poses that collapse is a unitary physical process. Asymptotic emergence employs algebraic quantum theory to demonstrate that some phenomena forbidden in the quantum limit (like collapse) are allowed in the classical limit due to some exponential sensitivity to perturbations in the Hamiltonian [118].
In a series of works [119, 120], Schonfeld has explored cloud chambers, Geiger counters, and Stern-Gerlach experiments for empirical tests of quantum measurement, including possible deviations from the Born rule in detecting rare events. He challenges the conventional view of measurement as an axiomatic feature of quantum mechanics, proposing that measurement emerges phenomenologically from collective interactions in quantum detection systems.
5.6.2 Mind-Inducing Collapse
A historically relevant but discredited hypothesis is that consciousness causes collapse. Wigner advocated for it until he learned about Zeh’s work on what would come to be known as decoherence [121]. We mainly mention this mind-matter duality here to disentangle a few ideas that are sometimes mixed up.
First, there is the actual proposal that consciousness causes the collapse, which presupposes a matter-mind duality. The discomfort this hypothesis raises in our physical understanding of the world has never been better embodied than in Bell’s quip [122], “Was the wave function of the world waiting to jump for thousands of millions of years until a single-celled living creature appeared? Or did it have to wait a little longer, for some better qualified system … with a PhD?”
Then, there is an epistemic proposal where the mind doesn’t objectively cause the collapse. Still, the collapse is just an illusion caused by the observer becoming aware of the measurement outcome. We find such perspective in London and Bauer [123, p. 251] who, in 1939, attributed the essential role played by consciousness to be “the increase of knowledge, acquired by the observation.” This is essentially the same hypothesis that sustains QBism nowadays, as discussed in Section 5.2.
Finally, Hameroff and Penrose [124] put forward precisely the opposite hypothesis: collapse causes consciousness. (And quantum coherences in the microtubules mentioned in Section 4 would be part of this process.) They argue that if consciousness cannot be reduced to an algorithmic process, then its physical origin must be in the stochastic nature of the wave function collapse, which, for Penrose, is an objective event, as discussed in Section 5.3.
5.7 Measurement in Quantum Field Theory
Quantum Field Theory (QFT) poses specific challenges to the Measurement Problem that are more daunting than in standard, non-relativistic quantum mechanics [125]. Because QFT integrates quantum mechanics and special relativity, locality is a built-in concept. This differs entirely from non-relativistic quantum mechanics, where locality is an add-on assumption, like when Bell imposed it in deriving his theorem. Sorkin showed that if one attempts to apply standard collapse theory in QFT, it may lead to superluminal information exchange in what he denoted impossible measurements [126, 127].
Figure 8 illustrates a version of Sorkin’s impossible measurement using our NH3 qubit. Alice, Bob, and Charlie are ready to perform measurements on a molecule-field coupled system at times (Alice), (Bob), and (Charlie). They are spaced so that causal effects are allowed between Bob and Alice and between Charlie and Bob but not between Alice and Charlie. In the relativity lingo, Alice and Charlie are spacelike separated.
Alice prepares a molecule in a state and can choose if she will flip it to at time or not. The molecule is coupled to an electromagnetic field. The field state is if the molecule is and if the molecule is . The field is initially and, after , it is updated according to Alice’s choice. At , Bob measures the field projected on . Charlie, who, at time , has no direct information about Alice’s choice, attempts to indirectly discover that by measuring the field at , which was updated after Bob’s intervention. Sorkins argues that, in principle, Charlie’s measurements on the field could reveal Alice’s choice, implying superluminal exchange. Therefore, avoiding such impossible measurements would require nontrivial constraints on which observables are consistent with causality.
Such imprecision in the description of the measurement in QFT has been characterized, with a flair to the dramatic, as “a major scandal in the foundations of quantum physics” [128]. Recently, Bostelmann, Fewster, and Ruep proposed that the situation is naturally solved within algebraic QFT as long as probes and couplings are constrained to be local [125, 129]. Their analysis is valid for general fields and spacetime geometries, making it a promising step toward a rigorous measurement theory in QFT. On the other hand, Polo-Gómez, Garay, and Martín-Martínez propose that a measurement theory in QFT could be built by exploring spatially smeared detector models [130].
6 The Ontological Status of the Wave function
One fundamental question underlying the Measurement Problem is whether the wave function represents an actual physical object (-Ontic) or is merely a mathematical tool encoding information (-Epistemic) [131]. The Many-Worlds Interpretation (Section 5.1), Objective Collapse Theories (Section 5.3), and Hidden-Variable Theories (Section 5.4) align with the -Ontic view, while epistemic interpretations (Section 5.2) are, by definition, -epistemic.
A common argument for the -Epistemic view is that the wave function evolves in an abstract, high-dimensional configuration space, making it unlikely to correspond to anything physically real [131]. However, such a criticism does not intimidate -ontic proponents, who argue that the wave function describes something fundamental about reality, even if not in three-dimensional space [132, 133]. Wave function realism, for example, holds that the wave function is a physical field existing in a high-dimensional configuration space [134]. Carroll and Singh take this high-dimensional -ontic stand to the extreme [41]. In an interpretation they call Mad-Dog Everettianism, they propose that the only fundamental elements of reality are the Hilbert Space, state vectors, and Hamiltonians. Everything else—including space, time, and classical variables—emerges from these.
Other -ontic interpretations have been developed to avoid the conceptual challenges of high-dimensional spaces. One approach is the multi-field interpretation, which states that an -body wave function is a muti-field that assigns properties to sets of points in the 3D space [135]. This is the view one of us adopted in an essay discussing the nature of the molecular wave function [136]. Gao [137] proposed an alternative -ontic interpretation in the 3D space that does not require multi-fields. In his view, the wave function represents a random discontinuous motion (RDM) of particles in real space. The wave function phase encodes momentum flow, giving it a concrete physical meaning. Moreover, RDM makes testable predictions, suggesting that collapse is an emergent effect of underlying stochastic motion rather than a fundamental postulate. Whether this model extends to relativistic Quantum Field Theory remains an open question.
Spacetime state realism reinterprets quantum mechanics by assigning quantum states to regions of spacetime rather than treating the wave function as an object in a high-dimensional space [138]. This approach aligns more naturally with Quantum Field Theory, which assigns quantum states to localized spacetime regions. A similar shift in emphasis to local quantities occurs in Density Functional Theory (DFT) [139], which, rather than working with the entire many-body wave function, pragmatically describes fermionic systems in terms of their density—a function of three spatial coordinates.
A different approach comes from Nomological Interpretations, which suggest that the wave function is not a physical object but a law of nature governing quantum behavior [132]. This is a common stance in the interpretation of Bohmian Mechanics [94, 140].
The Berry phase, which may arise during the cyclic evolution of quantum systems [141, 142], is a mathematical feature of the wave function that could be explored to assess its ontological status. Such a phase can build, for instance, when the nuclear geometry of a molecule evolves along a closed path encircling a conical intersection (a point of electronic state degeneracy). Unlike dynamic phases, the Berry phase is tied to the wave function’s geometric properties and may, in principle, manifest in observable interference effects. Valahu et al. [143] claimed to have observed such geometric-phase interference in a trapped-ion quantum simulator mimicking conical intersection dynamics, which would strongly support the -ontic interpretation. However, it’s unclear if such effects could be observed in actual molecules since their experiment could be considered more a controlled simulation than a molecular measurement. Moreover, Min et al. [144] and, more recently, Ibele et al. [145] have provided strong theoretical evidence that the Berry phase originates from the Born-Oppenheimer approximation and disappears in an exact electron-nuclear treatment, suggesting it may be an artifact rather than a fundamental quantum property.
Although the ontological nature of the wave function remains unsettled, recent experiments provided empirical support for wave function realism. Costello et al. [146] reconstructed Bloch wave functions in GaAs, demonstrating that wave functions are not just inferred but can be directly measured. Their work highlights that, at least in solid-state physics, the wave function has an observable physical structure. This challenges purely -epistemic interpretations and suggests that wave functions, at least in some contexts, correspond to real physical entities. Nevertheless, the scope of their findings is limited because the wave functions in the study describe quasiparticles (holes in semiconductors), they are not directly the same as the universal wave function of quantum mechanics.
7 Conclusion
It is easy to get lost in the Quantum Measurement Problem. The gargantuan amount of literature, the innumerable distinct theories, the interminable debates, and the entanglement between physics, philosophy, and formal logic all lend anyone stepping on this field a feeling analogous to being trapped in a dense tropical forest. This review intends to be an updated global map, which, although it cannot point out the escape route, as it simply doesn’t exist, will at least help us stay far from dangerous cliffs.
Here, we tried to offer a general overview of possible solutions to the Measurement Problem, which remains actively discussed by physicists and philosophers, and of newly proposed theories that we consider particularly promising or disruptive. They fall into five classes of explanations: Many-Worlds Interpretations, Epistemic Interpretations, Objective Collapse Theories, Hidden-Variable Theories, and Dualistic Collapse Hypotheses.
Today, the measurement problem is much less puzzling than it once was. Thanks to the decoherence program, we understand how quantum interferences are suppressed and how the environment selects the outcome basis. Nevertheless, we still lack a fundamental explanation for why only a single outcome is observed in experiments, and we are still taking the first steps toward a well-defined measurement theory in Quantum Field Theory. Indeed, given that QFT is likely the most successful physical theory ever developed, any satisfactory solution for the Measurement Problem must also be an acceptable solution for measurement in QFT. Therefore, it is worrisome to note that researchers on the Measurement Problem don’t usually deal with QFT (and vice versa).
Most scientists who use quantum mechanics in their daily work focus on its practical applications rather than its interpretation. Indeed, it makes no difference if we interpret the result of an experiment as the quantum state collapsing into a single branch or experiencing a branch that cuts communication from all others. Such subtle distinction belongs to defining a preferred worldview. It is even less relevant than choosing between Newtonian or Hamiltonian descriptions of classical dynamics, which have at least instrumental implications. The exception concerns objective collapse theories, which propose new physics beyond conventional quantum mechanics. In any case, it makes sense to keep challenging all interpretations and theories to prune the weak and reach complete consistency on the remaining, even if we never get a consensual single description of quantum mechanics.
Acknowledgments
This work received support from the French government under the France 2030 investment plan as part of the Initiative d’Excellence d’Aix-Marseille Université (A*MIDEX AMX-22-REAB-173 and AMX-22-IN1-48) and from the European Research Council (ERC) Advanced Grant SubNano (grant agreement 832237).